# Hybrid Beamforming with Widely-spaced-array for Multi-user Cross-Near-and-Far-Field Communications

### v1

@@ IV-B1 Alternating-Optimization-based Analog Beamforing in Sub-connected MU-WSA
Next, we fix $\mathbf{F}_{\mathrm{RF}}$ and extract the contribution of the analog combiner $\mathbf{W}_{\mathrm{RF}}$ to the sum SE (23). Since the spectral efficiency of each user is only related to its own combiner, we design the analog combiner for each user individually to maximize the overall spectral efficiency, which can be given as
@@
Algorithm 2 Hybrid beamforming with alternating optimization over submatrices in THz multi-user WSA systems
@@ IV-B2 Low-complexity Subarray-based Analog Beamforming in Fully-connected MU-WSA
Now we consider the fully-connected MU-WSA architecture where each subarray is connected to all the RF chains. To reduce the complexity of the analog beamforming, we propose a subarray-based steering vector reconstruction (SVR) scheme, which avoids the utilization of the high-complexity operation of SVD. Since the channel between each subarray and the users can be represented by the far-field PWM, we first simplify the analog beamforming design by designing the analog precoder for each subarray in parallel. Specifically, we aim to maximize the SE for each subarray, i.e.,
$SE_{k}\approx\log_{2}\operatorname{det}\left(\mathbf{I}+\sigma_{\mathrm{n}}^{-2}\mathbf{F}_{\mathrm{RF},k}^{H}\mathbf{H}_{k}^{H}\mathbf{W}_{\mathrm{RF}}\mathbf{W}_{\mathrm{RF}}^{H}\mathbf{H}_{k}\mathbf{F}_{\mathrm{RF},k}\right),$ (28)
where $\mathbf{H}_{k}$ is the PWM between the $k^{\mathrm{th}}$ subarray and all the users, and $\mathbf{F}_{\mathrm{RF},k}$ is the corresponding analog precoding matrix. Assuming a fixed analog combiner, then the analog precoder can be obtained through the beam steering vector. The steering vector between user $u$ and the first subarray can be represented as
@@
The distance between the $k^{\mathrm{th}}$ subarray and the user can be represented as
$D^{uk}=\sqrt{(x_{k}-x_{u})^{2}+(y_{k}-y_{u})^{2}+(z_{k}-z_{u})^{2}},$ (31)
where $p_{u}=(x_{u},y_{u},z_{u})$ and $p_{k}=(x_{k},y_{k},z_{k})$ are the position of user $u$ and the $k^{\mathrm{th}}$ subarray, respectively. Denote the distances between different reference antennas on the x- and z-axis as $d_{x}=\frac{(n_{x}-1)}{2}\lambda+d_{s}$ and $d_{z}=\frac{(n_{z}-1)}{2}\lambda+d_{s}$, where $n_{x}=n_{z}=\sqrt{\frac{N_{t}}{K}}$ is the number of antennas in one subarray at the x- and z-axis, and let $p_{1}=(0,0,z_{t})$, then we obtain $p_{k}=(x_{1}+(k_{x}-1)d_{x},0,z_{1}+(k_{z}-1)d_{z})=(x_{1}+\delta d_{x}^{k},0,z_{1}+\delta d_{z}^{k})$, where $k_{x}$ and $k_{z}$ are the x- and z-coordinate of subarray $k$. Therefore, applying the second order Taylor expansion that $\sqrt{1+\psi}\approx 1+\frac{1}{2}\psi-\frac{1}{8}\psi^{2}$, we can derive
@@
(33b) is obtained based on (16), and by considering equal subarray spacing at x- and z-axis, i.e., $d_{x}=d_{z}=d$. Therefore, the distance difference can be calculated as
@@ V Performance Evaluation
In this section, we compare the performance of the MU-WSA and the compact array for both 2-D and 3-D cases. We then evaluate the performance of the proposed algorithm in the MU-WSA architecture and compare it with the state-of-the-art algorithms in the compact architecture under different user scenarios. The simulation setup is summarized in Table I unless otherwise stated.
@@ V-A Cross Field Analysis of MU-WSA and Compact Array
We first compare the performance of MU-WSA and the compact array in 2-D scenarios where the transmitters and the receivers are at approximately the same height such that only the azimuth angle dominates. In Fig. LABEL:fig:2D-pattern, we analyze the beampattern of MU-WSA and compact array where four users are located at the same azimuth angle, ranging from $1$ to $20$m, represented by the white dot in the figure. Comparing Fig. LABEL:fig_2D:a and LABEL:fig_2D:b, we observe that the far-field analog beam in the compact array fails to distinguish these users since the beamforming gain along the target direction is the same with no distance domain differentiation. While in the MU-WSA, due to the near-field focusing effect, the analog beam pattern depicts the feature of additional distance domain resolution, providing the opportunity to mitigate interference for users with similar angles. Moreover, as the beam is more focused toward the target position rather than a target angle, the power can be more concentrated. To better illustrate this effect, in Fig. LABEL:fig:2D_SE, we show the performance of the sum SE versus the transmission distance. Here we neglect the influence of the beamforming algorithm and consider the upper bound of SE. Fig. LABEL:fig_2D_SE:a shows the sum SE when users are at the same azimuth angle. We observe that with increasing distance between the users and the BS, the MU-WSA architecture first demonstrates a large performance improvement compared to the traditional compact architecture. By contrast, since the compact array fails to cancel the inter-user interference at the same angle, the sum SE drops to approximately zero because of the low SINR. When the communication distance increases to a certain point, the near-field beam-focusing effect vanishes, resulting in a smaller gap between the performance of the MU-WSA and the compact array. On the other hand, for users at different azimuth angles as shown in Fig. LABEL:fig_2D_SE:b, both the MU-WSA and the compact array architecture are able to distinguish the users and the sum SE is almost the same at all distances.
@@
In this section, we consider the 3D scenarios where both the azimuth angle $\phi$ and the elevation angle $\theta$ should be considered. Therefore, even when the users are at the same azimuth angle, the different elevation angles $\phi_{u}$ can support transmission to these users. However, when the distance between users and the BS becomes larger, the elevation angles of different users become approximately the same, i.e., $\phi_{u}\approx\phi,\forall u$. In this case, as shown in Fig. LABEL:fig:3D-pattern, the beamforming gain of the compact and the MU-WSA shows a similar behavior as the 2D case where the MU-WSA provides additional distance domain resolution and depicts the beam focusing effect.
@@
Channel model CNFF PWM
@@
Capable of transmitting number of data streams larger than number of multipath components Yes No
@@ V-B Performance Analysis of Hybrid Beamforming Algorithms
In Fig. 11, we first evaluate the sum SE of the proposed algorithms versus the transmit power. The users are distributed randomly within a sector of $120^{\circ}$ angle and a radius ranging from $1$ to $20$m. We compare the performance with the optimal upper bound and an SVD-based benchmark algorithm HySBD-2 proposed in [10]. From the figure, we observe that the MU-WSA architecture always outperforms the compact array regardless of the algorithms used. Specifically, comparing the upper bound of the two architectures, the MU-WSA improves the SE by over $60$% at a power of $20$dBm. In the FC system, the proposed SVR algorithm can achieve over 95% sum SE of the upper bound and the existing SVD-based algorithm, but with much lower computational complexity. Moreover, the proposed AO algorithm in SC achieves over 80% of the sum SE in the FC system with a lower power consumption as the SC system reduces the number of phase shifters by $K^{2}$.
@@
In Fig. 13, we evaluate the sum SE versus the number of users. We observe that, when the number of users is small, namely, users are sparsely distributed in space, the performance of the WSA and the compact array are similar. However, as the number of users becomes larger to a certain point, the performance of the compact array system experiences a rapid drop due to the similar angles of the users.
In Table III, we analyze the computational complexity and the running time of the proposed algorithms. For SVR which utilizes a predefined steering vector, the precoder at each subarray for each user can be calculated through through simple calculation, resulting in the lowest complexity, i.e., $O(UK)$. While the benchmark algorithm HySBD-2 [10] conducts SVD operation on the channel matrix of all users, resulting in a complexity of $O((UL_{r})^{2}N_{t})$. These two algorithms can be applied to the FC system while the proposed steering vector-based scheme has a much lower complexity. For the more complicated SC beamforming scheme, the proposed AO algorithm is of the highest complexity since it requires the calculation of the inverse of the matrix $\mathbf{B}_{j}\in\mathbb{C}^{\frac{(K-1)L_{t}}{K}\times\frac{(K-1)L_{t}}{K}}$ in (25b), which has a complexity of $O((K-1)^{3}L_{t}^{3}/K^{3})$, and the SVD operation for $\mathbf{X}_{j}\in\mathbb{C}^{\frac{N_{t}}{K}\times\frac{N_{t}}{K}}$, which has a complexity of $O((\frac{N_{t}}{K})^{3})$. Therefore, the overall complexity after $K$ iterations is $O(((K-1)^{3}L_{t}^{3}+N_{t}^{3})/K^{2})$. The running time of the algorithms with MatLab R2022a is also summarized to directly demonstrate the complexity. We observe that the running time corresponds to the complexity analysis, where SVR has the lowest complexity. While AO has the highest complexity but is able to solve the more complex SC problem with acceptable performance loss.
@@ VI Conclusion
AO (SC) $O(((K-1)^{3}L_{t}^{3}+N_{t}^{3})/K^{2})$ around 0.35s
In this paper, we extended the WSA hybrid beamforming architecture into the multi-user scenarios, which can utilize both the angular and distance domain resolutions. Since PWM cannot be applied in the enlarged near-field region of MU-WSA, we investigated the CNFF channel model. Then, by considering both SC and FC systems, we formulated a joint optimization problem to maximize the sum SE. We then decomposed the problem into the design of the MU-WSA architecture, i.e., the subarray spacing $d_{s}$ and the number of subarrays $K$, and the hybrid beamformers. Firstly, based on the analysis of the $(K.d_{s})$ pair, we proposed a search-based algorithm to find the optimal solution for the architecture design. Then, for SC systems where the analog precoder has a block-diagonal format, we propose the AO algorithm to iteratively optimize each submatrix in the analog precoder. In the FC system, a low-complexity SVR algorithm is proposed where the analog beamformers are found through the reconstruction of the far-field steering vectors. We evaluate the beampattern of MU-WSA and the compact array in 2D and 3D cases. We show that in 2D cases, the MU-WSA outperforms the traditional compact array when users are located at the same azimuth angle. While in 3D scenarios, the compact array and the MU-WSA exhibit comparable performance at short distances, but as the distance increases, MU-WSA depicts a gradual superiority over the compact array. When users are located at different angles, the two architectures can achieve similar performance. Therefore, to conclude, the MU-WSA architecture is preferred when the users are more likely to locate at similar angles or directions, while both the MU-WSA and the compact array are feasible when users tend to be scattered in the angle domain. Furthermore, we compare the performance of the proposed AO and SVR algorithms. We show that the proposed AO algorithm in the SC system can achieve over 80% of the sum SE in the FC system but with lower hardware complexity. Moreover, the proposed SVR achieves over 95% of the upper bound of SE, but much lower computational complexity.
@@
where $(i_{1},j_{1})$ and $(i_{2},j_{2})$ are the x- and z-coordinate of antenna $i$ and $j$, respectively. Define $h(d_{s})=\frac{1}{\pi}(\beta_{k,i,j}-\beta_{m,i,j})$, we next prove $h(d_{s})<1$. Specifically, we can derive
$h(d_{s})=(\beta_{k,i,j}-\beta_{m,i,j})/\pi$ (39)
$=(j_{1}-i_{1})[(x_{u}-x_{k})(\frac{1}{D^{uk}}-\frac{1}{D^{um}})+(m_{x}-k_{x})\frac{d_{s}}{D^{um}}]$
$+(j_{2}-i_{2})[(z_{u}-z_{k})(\frac{1}{D^{uk}}-\frac{1}{D^{um}})+(m_{z}-k_{z})\frac{d_{s}}{D^{um}}]$
$\leq(\sqrt{Nr}-1)[\left|x_{u}-x_{k}\right|\cdot\left|\frac{1}{D^{uk}}-\frac{1}{D^{um}}\right|+(\sqrt{K}-1)\frac{d_{s}}{D^{um}}]$
$+(\sqrt{Nr}-1)[\left|z_{u}-z_{k}\right|\cdot\left|\frac{1}{D^{uk}}-\frac{1}{D^{um}}\right|+(\sqrt{K}-1)\frac{d_{s}}{D^{um}}]$
@@
$\quad\quad\quad\quad\quad\quad\left.+2(\sqrt{K}-1)\frac{d_{s}}{D^{um}}\right]$
@@
$\begin{split}h(d_{s})&\leq(\sqrt{Nr}-1)\left[\frac{(\left|x_{u}-x_{k}\right|+\left|z_{u}-z_{k}\right|)}{D^{uk}}\frac{S_{t}}{D^{uk}+S_{t}}\right.\\ &\quad\quad\quad\quad\quad\quad\left.+2(\sqrt{K}-1)\frac{d_{s}}{D^{um}}\right]\end{split}$ (40a)
$\leq(\sqrt{Nr}-1)(\frac{\sqrt{2}S_{t}}{D^{um}}+2(\sqrt{K}-1)\frac{d_{s}}{D^{um}}),$ (40b)

### v2

@@ IV-B1 Alternating-Optimization-based Analog Beamforming in Sub-connected MU-WSA
Next, we fix $\mathbf{F}_{\mathrm{RF}}$ and extract the contribution of the analog combiner $\mathbf{W}_{\mathrm{RF}}$ to the sum SE (28a). Since the spectral efficiency of each user is only related to its own combiner, we design the analog combiner for each user individually to maximize the overall spectral efficiency, which can be given as
@@
Algorithm 1 Hybrid beamforming with subarray-based alternating optimization (S-AO) in THz multi-user WSA systems
@@ IV-B2 Low-complexity Subarray-based Analog Beamforming in Fully-connected MU-WSA
Now we consider the fully-connected MU-WSA architecture where each subarray is connected to all the RF chains. In this case, we assume each user is equipped with one RF chain to receive $N_{s}=1$ data stream. Then, to reduce the complexity of the analog beamforming, we propose a subarray-based steering vector reconstruction (SVR) scheme, which avoids the utilization of the high-complexity operation of SVD. Since the channel between each subarray and the users can be represented by the far-field PWM, we first simplify the analog beamforming design by designing the analog precoder for each subarray. We denote the overall channel matrix as $\mathbf{H}=[\mathbf{H}_{sub}^{1},\cdots,\mathbf{H}_{sub}^{K}]$, where $\mathbf{H}_{sub}^{k}$ represents the channel matrix between the $k^{\mathrm{th}}$ subarray and all the users. Then, based on eq. (28b), we have
$SE$ $\approx\sum_{k=1}^{K}\log_{2}\operatorname{det}\left(\mathbf{I}+\sigma_{\mathrm{n}}^{-2}\mathbf{F}_{\mathrm{RF},k}^{H}(\mathbf{H}_{sub}^{k})^{H}\mathbf{W}_{\mathrm{RF}}\right.$ (33)
×$\left.\quad\quad\quad\quad\times\mathbf{W}_{\mathrm{RF}}^{H}\mathbf{H}_{sub}^{k}\mathbf{F}_{\mathrm{RF},k}\right)$
$=\sum_{k=1}^{K}SE_{k}$
Since $SE_{k}$ is independent of each other assuming fixed analog combiner, we can design the precoder $\mathbf{F}_{\mathrm{RF},k}$ in parallel. Specifically, the analog precoder can be obtained through the beam steering vector where the steering vector between user $u$ and the first subarray is represented as
@@
Denote $\delta d_{x}^{k}=(k_{x}-1)d$ and $\delta d_{z}^{k}=(k_{z}-1)d$. Then, based on equation (9), and applying the second order Taylor expansion that $\sqrt{1+\psi}\approx 1+\frac{1}{2}\psi-\frac{1}{8}\psi^{2}$, we can derive
@@
Therefore, the distance difference can be calculated as
@@ V Performance Evaluation
In this section, we compare the performance of the MU-WSA and the compact array for both 2-D and 3-D cases. We then evaluate the performance of the proposed algorithm in the MU-WSA architecture and compare it with the state-of-the-art algorithms in the compact architecture under different user scenarios. The simulation setup is summarized in Table I where we set the maximum array aperture as $S_{t}^{max}=1$ m, and fix the number of subarrays as $K=4$. Then the subarray spacing $d_{s}=0.692$ m can be obtained through eq. (24).
@@ V-A Cross Field Analysis of MU-WSA and Compact Array
We first compare the performance of MU-WSA and the compact array in 2-D scenarios where the transmitters and the receivers are at approximately the same height such that only the azimuth angle dominates. In Fig. LABEL:fig:2D-pattern, we analyze the analog beampattern of MU-WSA and compact array where four users are located at the same azimuth angle, ranging from $1$ to $20$m, represented by the white dots shown in the figure. Comparing Fig. LABEL:fig_2D:a and LABEL:fig_2D:b, we observe that the far-field analog beam in the compact array fails to distinguish these users since the beamforming gain along the target direction is the same with no distance domain differentiation. While in the MU-WSA, due to the near-field focusing effect, the analog beam pattern depicts the feature of additional distance domain resolution, providing the opportunity to mitigate interference for users with similar angles. Moreover, as the beam is more focused toward the target position rather than a target angle, the power can be more concentrated. Additionally, we observe the existence of grating lobe from Fig. LABEL:fig_2D:b which is caused by the widely-spaced subarrays. The grating lobe would introduce additional user interference into the system, which may require interference cancellation beamforming algorithms, such as the baseband BD algorithm depicted in Sec. IV-A.
In Fig. LABEL:fig:2D_SE, we evaluate the performance of the sum SE versus the transmission distance. To focus on the performance of the antenna array architecture, we temporarily neglect the effect of different connection types and the analog beamforming algorithms, and only consider the fully-digital scenario where the upper bound of the sum SE is achieved by applying the digital BD algorithm [10, 9]. Fig. LABEL:fig_2D_SE:a shows the sum SE when users are at the same azimuth angle. We observe that with increasing distance between the users and the BS, the MU-WSA architecture first demonstrates a large performance improvement compared to the traditional compact architecture. By contrast, since the compact array fails to cancel the inter-user interference at the same angle, the sum SE drops to approximately zero because of the low SINR. When the communication distance increases to a certain point, the near-field beam-focusing effect vanishes, resulting in a smaller gap between the performance of the MU-WSA and the compact array. On the other hand, for users at different azimuth angles as shown in Fig. LABEL:fig_2D_SE:b, both the MU-WSA and the compact array architecture are able to distinguish the users. However, due to the additional inter-user interference caused by the grating lobe in WSA, the SE in the compact array is slightly larger than that in the WSA array. Fortunately, due to the interference-cancellation-based BD algorithm, the difference is negligible across varying distances.
@@
In this section, we consider the 3D scenarios where both the azimuth angle $\phi$ and the elevation angle $\theta$ should be considered. Therefore, even when the users are at the same azimuth angle, the different elevation angles $\phi_{u}$ can support transmission to these users. However, when the distance between users and the BS becomes larger, the elevation angles of different users become approximately the same, i.e., $\phi_{u}\approx\phi,\forall u$. In this case, as shown in Fig. LABEL:fig:3D-pattern, the beamforming gain of the compact and the MU-WSA show a similar behavior as the 2D case where the MU-WSA provides additional distance domain resolution and depicts the beam focusing pattern, at the cost of the grating lobe effect.
@@
Channel model CNFF or SWM PWM
@@
Capable of transmitting more data streams than number of multipath components Yes No
@@ V-B Performance Analysis of Hybrid Beamforming Algorithms
In Fig. 11, we first evaluate the sum SE of the proposed algorithms versus the transmit power. The users are distributed randomly within a sector of $120^{\circ}$ angle and a radius ranging from $1$ to $20$m. We compare the performance with the optimal upper bound and an SVD-based benchmark algorithm HySBD-2 proposed in [10]. From the figure, we observe that the MU-WSA architecture always outperforms the compact array regardless of the algorithms used. Specifically, comparing the upper bound of the two architectures, the MU-WSA improves the SE by over $60$% at a power of $20$dBm. In the FC system, the proposed SVR algorithm can achieve over 95% sum SE of the upper bound and the existing SVD-based algorithm, but with much lower computational complexity. Moreover, the proposed S-AO algorithm in SC achieves over 80% of the sum SE in the FC system with a lower power consumption as the SC system reduces the number of phase shifters by $K^{2}$.
@@
In Fig. 13, we evaluate the sum SE versus the number of users. We observe that, when the number of users is small, i.e., the users are sparsely distributed in space, the performance of the MU-WSA and the compact array are similar. However, as the number of users increases, the compact array system experiences a rapid performance degradation. This is because, as the user density increases, users are more likely to be located at similar angles. Since the compact array relies on the angular resolution of far-field communication, it fails to mitigate interference from users at similar angles, thereby reducing the sum SE. In contrast, the MU-WSA system leverages the near-field distance domain, which enables efficient interference suppression even among users with identical angles. As a result, the WSA system can effectively serve all the users simultaneously, maintaining higher spectral efficiency.
In Table III, we analyze the computational complexity and the running time of the proposed algorithms. For SVR which utilizes a predefined steering vector, the precoder at each subarray for each user can be calculated through through simple calculation, resulting in the lowest complexity, i.e., $O(UK)$. While the benchmark algorithm HySBD-2 [10] conducts SVD operation on the channel matrix of all users, resulting in a complexity of $O((UL_{r})^{2}N_{t})$. These two algorithms can be applied to the FC system while the proposed steering vector-based scheme has a much lower complexity. For the more complicated SC beamforming scheme, the proposed S-AO algorithm is of the highest complexity since it requires the calculation of the inverse of the matrix $\mathbf{B}_{j}\in\mathbb{C}^{\frac{(K-1)L_{t}}{K}\times\frac{(K-1)L_{t}}{K}}$ in (30b), which has a complexity of $O((K-1)^{3}L_{t}^{3}/K^{3})$, and the SVD operation for $\mathbf{X}_{j}\in\mathbb{C}^{\frac{N_{t}}{K}\times\frac{N_{t}}{K}}$, which has a complexity of $O((\frac{N_{t}}{K})^{3})$. Therefore, the overall complexity after $K$ iterations is $O(((K-1)^{3}L_{t}^{3}+N_{t}^{3})/K^{2})$. The running time of the algorithms with MatLab R2022a is also summarized to directly demonstrate the complexity. We observe that the running time corresponds to the complexity analysis, where SVR has the lowest complexity. While S-AO has the highest complexity but is able to solve the more complex SC problem with acceptable performance loss.
@@ VI Conclusion
S-AO (SC) $O(((K-1)^{3}L_{t}^{3}+N_{t}^{3})/K^{2})$ around 0.35s
In this paper, we extended the WSA hybrid beamforming architecture into the multi-user scenarios, which can utilize both the angular and distance domain resolutions. Since PWM cannot be applied in the enlarged near-field region of MU-WSA, we investigated the CNFF channel model. Then, by considering both SC and FC systems, we formulated a joint optimization problem to maximize the sum SE. We then decomposed the problem into the design of the MU-WSA architecture, i.e., the subarray spacing $d_{s}$ and the number of subarrays $K$, and the hybrid beamformers. Firstly, based on the analysis of the $(K.d_{s})$ pair, we derive the optimal subarray spacing for the architecture design. Then, for SC systems where the analog precoder has a block-diagonal format, we propose the S-AO algorithm to iteratively optimize each submatrix in the analog precoder. In the FC system, a low-complexity SVR algorithm is proposed where the analog beamformers are found through the reconstruction of the far-field steering vectors. We evaluate the beam pattern of MU-WSA and the compact array in 2D and 3D cases. We show that in 2D cases, the MU-WSA outperforms the traditional compact array when users are located at the same azimuth angle. While in 3D scenarios, the compact array and the MU-WSA exhibit comparable performance at short distances, but as the distance increases, MU-WSA depicts a gradual superiority over the compact array. When users are located at different angles, the two architectures can achieve similar performance. Therefore, to conclude, the MU-WSA architecture is preferred when the users are more likely to locate at similar angles or directions, while both the MU-WSA and the compact array are feasible when users tend to be scattered in the angle domain. Furthermore, we compare the performance of the proposed S-AO and SVR algorithms. We show that the proposed S-AO algorithm in the SC system can achieve over 80% of the sum SE in the FC system but with lower hardware complexity. Moreover, the proposed SVR achieves over 95% of the upper bound of SE, but much lower computational complexity.
@@
where $(i_{1},j_{1})$ and $(i_{2},j_{2})$ are the x- and z-coordinate of antenna $i$ and $j$, respectively. We denote $h(d)=\frac{1}{\pi}|\beta_{k,i,j}-\beta_{m,i,j}|$ where $d=\frac{(n-1)}{2}\lambda+d_{s}$. Next, we prove $h(d)<1$. Specifically, we can derive
$h(d)=|\beta_{k,i,j}-\beta_{m,i,j})|/\pi$
$=|(j_{1}-i_{1})[(x_{u}-x_{k})(\frac{1}{D^{uk}}-\frac{1}{D^{um}})+(m_{x}-k_{x})\frac{d}{D^{um}}]$
$+(j_{2}-i_{2})[(z_{u}-z_{k})(\frac{1}{D^{uk}}-\frac{1}{D^{um}})+(m_{z}-k_{z})\frac{d}{D^{um}}]|$
$\leq|(j_{1}-i_{1})[(x_{u}-x_{k})(\frac{1}{D^{uk}}-\frac{1}{D^{um}})+(m_{x}-k_{x})\frac{d}{D^{um}}]|$
$+|(j_{2}-i_{2})[(z_{u}-z_{k})(\frac{1}{D^{uk}}-\frac{1}{D^{um}})+(m_{z}-k_{z})\frac{d}{D^{um}}]|$
$\leq(\sqrt{Nr}-1)[\left|x_{u}-x_{k}\right|\cdot\left|\frac{1}{D^{uk}}-\frac{1}{D^{um}}\right|+(\sqrt{K}-1)\frac{d}{D^{um}}]$
$+(\sqrt{Nr}-1)[\left|z_{u}-z_{k}\right|\cdot\left|\frac{1}{D^{uk}}-\frac{1}{D^{um}}\right|+(\sqrt{K}-1)\frac{d}{D^{um}}]$
@@
$\quad\quad\quad\quad\quad\quad\left.+2(\sqrt{K}-1)\frac{d}{D^{um}}\right]$ (43)
@@
$\begin{split}h(d)&\leq(\sqrt{Nr}-1)\left[\frac{(\left|x_{u}-x_{k}\right|+\left|z_{u}-z_{k}\right|)}{D^{uk}}\frac{S_{t}}{D^{uk}+S_{t}}\right.\\ &\quad\quad\quad\quad\quad\quad\left.+2(\sqrt{K}-1)\frac{d}{D^{um}}\right]\end{split}$ (44a)
$\leq(\sqrt{Nr}-1)(\frac{\sqrt{2}S_{t}}{D^{um}}+2(\sqrt{K}-1)\frac{d}{D^{um}}),$ (44b)